The sum of two integers is sixteen. Five tiems the smaller integer is eight more than three times the larger integer. Find the larger integer.
Answer by lwsshak3(6484)

The sum of two integers is sixteen. Five tiems the smaller integer is eight more than three times the larger integer. Find the larger integer.
**
let x=larger integer
16-x=smaller integer
..
5(16-x)=3x+8
80-5x=3x+8
8x=72
x=9